OpenStudy (anonymous):

determine the max and min values of f(x)=2sinx+sin(2x) on the interval [0,2pi]

OpenStudy (turingtest):

step 1: take the derivative

OpenStudy (anonymous):

I don't know if @mariaad is here. But I guessed at the end. I wonder if there is a simple way to solve that explicitly? Or is it double angle and all that...

OpenStudy (anonymous):

i am here sorry!

OpenStudy (anonymous):

for the derivative i got y= 2cosx + 2cos(2x)

OpenStudy (anonymous):

i just don't know when i set it equal to 0 what to do

OpenStudy (turingtest):

yeah it's easy with \[\cos(2x)=2\cos^2x-1\]so using that it will be quadratic in cosx

OpenStudy (anonymous):

i dont understand

OpenStudy (rogue):

You got y = 2cos(x) + 2cos(2x) You can make the substitution cos(2x) = 2cos^2 (x) - 1 So you have y = 2cos(x) + 2(2cos^2 (x) - 1) y = 4 cos^2 x + 2 cos(x) - 2 Which you can solve like a quadratic equation.